Y=x^2+8x-2 in vertex form
Answer by Theo(13342)

here's a reference.

https://mathbitsnotebook.com/Algebra1/Quadratics/QDVertexForm.html

you are basically completing the squares as follows:

start with y = x^2 + 8x - 2

add 2 to both sides of the equation to get y + 2 = x^2 + 8x

complete the square on the expression on the right side of the equation to get:

y + 2 = (x+4)^2 - 16

subtract 2 from both sides of the equation to get y = (x+4)^2 - 18

that's your vertex form.

there's another way.

standard form of the equation is y = ax^2 + bx + c

a is the coefficient of the x^2 term.
b is the coefficient of the x term.
c is the constant term.

your equation is y = x^2 + 8x - 2

therefore you get a = 1 and b = 8 and c = -2

the standard form of the vertex form of the equation is y = a * (x-h)^2 + k

h is equal to -b/2a = -8/2 = -4

(x-h) is therefore equal to (x - (-4)) = (x+4)

k is equal to f(h) = (-4)^2 + 8 * (-4) - 2 which is equal to 16 - 32 - 2 which is equal to -16 - 2 which is equal to -18.

a is equal to 1, therefore y = a * (x-h)^2 + k becomes y = 1 * (x+4)^2 - 18 which becomes y = (x+4)^2 - 18.

that's your solution.